OpenStudy (dancergirl45):

Which of the following options is an equivalent function to f(x) = 2(5)^2x? f(x)=50^x f(x)=100^x f(x)=2(25)^x f(x)=4(25)^x

OpenStudy (dancergirl45):

@YoungStudier @welshfella

OpenStudy (dancergirl45):

@AloneS

OpenStudy (solomonzelman):

For example: \(\large\color{black}{ \displaystyle 3^{4x} =\left(3^4\right)^x=81^x }\) and therefore, if I had: \(\large\color{black}{ \displaystyle 6\times 3^{4x} =6\times \left(3^4\right)^x=6\times 81^x }\)

OpenStudy (solomonzelman):

You are using the same rules, but with different numbers. Hope this helped.

OpenStudy (dancergirl45):

so then the answer is D

OpenStudy (solomonzelman):

did the coefficient "6", in my example change ?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 2\times 5^{2x} =2\times \left(5^2\right)^x=\quad? }\)

OpenStudy (dancergirl45):

25

OpenStudy (dancergirl45):

50

OpenStudy (dancergirl45):

are you still there?

OpenStudy (solomonzelman):

Yes, but I have a feeling that you need more RVW on the topic.

OpenStudy (dancergirl45):

well can u help me

OpenStudy (solomonzelman):

Lets try this once again

OpenStudy (solomonzelman):

\(\large\color{blue}{ \displaystyle f(x) = 2(5)^{2x} }\) \(\large\color{blue}{ \displaystyle f(x) = 2(5^2)^{x} }\) \(\large\color{black}{ \displaystyle 5^2=5\cdot 5 = 25 }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(x) = ? }\)

OpenStudy (dancergirl45):

25

OpenStudy (solomonzelman):

You mean that f(x)=25 ??

OpenStudy (dancergirl45):

yes

OpenStudy (solomonzelman):

you have dropped the exponent and the coefficient....

OpenStudy (solomonzelman):

you seem to be too behind...